The current of 2.0 amperes passes through a cell of e.m.f. 1.5 volts having internal resistance of 0.15? . The potential difference measured in volts across both the terminals of the cell will be
A. 1.35 B. 1.50 C. 1.00 D. 1.20

Grade:

2 Answers

Arun
3 years ago
ANSWER
Given that,
Current = 0.2 A
Internal resistance = 0.15 ohm
e. m. f = 1.5 volt
We know that,
  V=E−ir
 V=1.5−0.2×0.15
 V=1.5−0.03
 V=1.47volt
Hence, the potential difference is 1.47 V across both the ends of the cell
Vikas TU
3 years ago
Terminal voltage is the voltage output of a device is measured across its terminals. Terminal voltage is calculated by V = emf – Ir.